In an electric lamp, 2.50C pass through each second. How many electrons does this represent?


1
Expert's answer
2022-03-21T12:42:37-0400

The number of electrons passing through the lamp per second

"N=\frac{q}{e}=\frac{2.50}{1.60\times 10^{-19}}=1.56\times 10^{19}"